### v1

@@ Abstract
Cyber-Physical Systems (CPSs) tightly integrate computation with physical entities, often generating vast amounts of time series data from thousands of sensors. Although knowledge graphs offer a powerful means to contextualize these data, existing approaches to integrating knowledge graphs with time series data lack a concept to model the continuous temporal values inherent in CPSs. This gap can make expressing computations on the sensor data cumbersome. In this work, we propose the integration of knowledge graphs and signals, a proven concept for modeling temporal values. By treating signals as first-class citizens in query languages, we can enable seamless querying over knowledge graphs and signals. While the knowledge graph captures information on the CPS, signals represent its run-time data from sensors. We discuss the implications of such an approach and propose SigSPARQL, an extension to the SPARQL query language, to demonstrate these concepts. Furthermore, we evaluate the feasibility of implementing SigSPARQL with a prototype and demonstrate the applicability of the query language for a monitoring use case within a CPS.
@@
However, fully utilizing sensor data in CPS is challenging. Firstly, data are frequently siloed across various systems. When this occurs, it is essential to have unified data access to effectively combine information from these sources. In addition, even if sensor data access does not pose a problem, some tasks require contextualizing the sensor data. In other words, the tasks require information on the CPS itself, not just the pure sensor data (e.g., which power consumers exist). For both aspects, data access and contextualization, Semantic Web technologies can provide a solution.
Materializing sensor data in a graph introduces additional overhead [36]. Other systems, such as time series databases [23], specialize in storing sensor data and can therefore avoid this overhead. However, such systems often do not provide advanced contextualization capabilities. This has led to Ontology-Based Data Access (OBDA) [32] approaches in CPSs (e.g., [35, 6, 25]) that provide a virtual knowledge graph on top of specialized data stores using an appropriate mapping. The resulting graph can be queried in conjunction with the context information that is stored in a regular Resource Description Framework (RDF) store [2], thus enabling unified and contextualized data access.
Although OBDA addresses many problems in the context of CPSs, formulating computations on time series data can become cumbersome. After mapping a time series to an RDF graph, the value of a single sensor can be distributed over thousands of elements that represent individual observations. This can lead to an increase in complexity, as queries must reconstruct the concept of a temporal value from the individual observations.
While efforts have been made to integrate knowledge graphs and time series data (e.g., [8]), they do lack a concept that abstracts over individual observations to alleviate this complexity. As a result, users must resort to “workarounds” to express computations such as summing up the temporal values from two sensors. Although some requirements can be met with these workarounds, there is a mismatch between the conceptual nature of the computation and the formulation in a query language. Signals [22] is a concept that can abstract over individual observations.
This work proposes to support signals as “first-class citizens” when querying knowledge graphs to overcome the shortcomings of contemporary approaches when formulating computations over time series data. The integration of this concept involves rethinking the semantics of queries and their results to accommodate the temporal dimension of signals. When considering this integration, the following research questions emerge.
@@
This work is structured as follows. Section 2 discusses related work, while Section 3 presents a motivating example. Then, Section 4 provides a background on signals, and Section 5 discusses their relationship to time series data. Section 6 presents a list of requirements to bridge the gap between knowledge graphs and signals and proposes a set of operators that address these requirements. Next, Section 7 applies these concepts to the SPARQL query language. Section 9 applies the proposed SPARQL extension to the motivating example, while Section 8 assesses its feasibility with a prototype. Finally, Section 10 reflects on the proposed approach, while Section 11 concludes this work.
@@ 2 Related Work
We have identified four key areas related to our research topics: (i) Querying temporal data in RDF streams, (ii) Querying temporal data with OBDA, (iii) Combining continuous querying and OBDA, and (iv) Integrating graphs with time series data.
Querying temporal data in RDF streams: The main objective of RDF Stream Processing (RSP) is to enable continuous queries on streaming graph data. As this technology has been used for monitoring CPSs (e.g., [19]), it is relevant for this discussion. Research in RSP can be broadly categorized into two main approaches: window-based systems and those inspired by Complex Event Processing (CEP).
On the one hand, window-based approaches (e.g., [9, 7, 26, 16]), focus on querying streaming data by dividing unbounded streams into bounded relations using window operators. Rost et al. [33] introduced a similar approach to property graphs. TEF-SPARQL [18] extends this paradigm by incorporating facts into the query language. Using facts, one can remember assertions beyond a single window.
On the other hand, CEP-inspired RSP systems (e.g., [4, 15]) aim to detect patterns in event streams and aggregate them into high-level events. These efforts integrate CEP concepts into continuous query frameworks based on SPARQL. OntoEvent [28] proposes a different approach by formulating CEP queries within an ontology-based knowledge base. Furthermore, PipeFlow [34], proposes embedding the queries in a data flow programming language, while Teymourian et al. [37] embed SPARQL queries in a logic programming environment.
Querying temporal data with OBDA: The emergence of the OBDA paradigm, marked by approaches such as Mastro [13] and Ontop [12], has increased the applicability and adoption of RDF-based methods. This success stems from the concept of a virtual knowledge graph layer built on top of existing databases. Although the initial focus of these approaches has primarily been on relational databases, recent approaches focus on building virtual knowledge graphs for time series data. Steindl et al. [35] demonstrated an approach to integrate time series data from OPC UA, a widely adopted standard in the manufacturing domain. In addition, Chrontext [6] proposes an approach for rewriting SPARQL queries to queries on time series databases. Furthermore, Ontop-temporal [24] extends the querying capabilities by allowing users to evaluate temporal logic formulae over the knowledge graph.
@@
Approaches integrating graphs with time series data: Recently, the combination of graph and time series data has received increased attention. Bollen et al. [8] present an approach that extends matching graphs with measurement and time series patterns. The bindings of these patterns can then be used to access the underlying time series. Furthermore, Ammar et al. [3] outline a research roadmap with the goal of creating a hybrid data model for graphs and time series data.
The approach proposed in this work can also be attributed to the latter category. Our approach differentiates itself from the mentioned works by relying on signals [22] to model temporal values and treating time implicitly. Using signals allows us to build on existing work created in the Functional Reactive Programming (FRP) community and tackle problems that arise from using an observation-based approach (see Section 5).
@@ 3 A Motivating Example
This section introduces a motivating example of monitoring a CPS that is used throughout this work. Figure 1 shows a conceptual system model of a garage that contains multiple Electric Vehicle (EV) chargers that the building occupants are free to use and a battery to mitigate peaks in overall power consumption. However, the garage is a substantial power consumer in the area. Therefore, without any further restriction, supporting the garage’s peak power demands would require additional power grid infrastructure, resulting in additional costs for the involved parties.
An envisioned approach to circumvent additional infrastructure is to impose an operating envelope, limiting the maximum power the garage can drain from the grid. This limitation places the responsibility of respecting this threshold on the operator of the EV chargers, who may be held liable for any violations. Thus, the garage operator needs to be able to detect envelope violations to present alerts to the responsible employees.
@@
Figure 2 depicts a conceptual model of the charging garage expressed in an extended data model proposed. The data model is based on RDF. In addition to a regular RDF graph (white ovals and solid lines), the extended data model introduces signals (green circles and dashed lines). For now, it suffices to know that the signals can model temporal values and that the dashed connections can associate multiple signals with a single node in the graph. These definitions will be detailed in Section 7.1.
In this example, the <Battery1> has two signals. models the active power, while models the battery’s State of Charge (SoC). In addition to the battery, other devices within the garage have an active power signal (e.g., <Charger1>). Lastly, the garage itself has a signal that models the operating envelope. As there is no active power signal at the garage level, we must compute this signal based on the active power of the contained entities (e.g., chargers). Note that the set of contained devices is not known at design-time and must be extracted by querying the knowledge graph. A query that implements this requirement will be presented in Section 9.
@@ 4 Functional Reactive Programming
Assume that the expression SineWave() in Listing 1 creates a signal representing a sine wave. Note that the value of a implicitly depends on time. Therefore, evaluating a at two different time points may yield different results. As a is a signal, the expression b = a * 2 denotes that b is twice the value of a at any given time point. This is different from denoting that b is twice the current value of a in imperative programming languages. In FRP, if a changes its value, the value of b automatically reflects this change, allowing b and c to become scaled sine waves. The concept of a signal is very flexible and allows us to model sensor values over time.
Table 1 contains descriptions of relevant FRP concepts and relates these concepts to the motivating example. Formally, a signal (line 1) is a partial function of a time domain to a value domain . The value of a signal at a time point is denoted by . We refer to any function that has signals as input or output as signal function111This is a more lenient usage than the original definition [22], which restricts signal functions to achieve favorable computational properties. In this work, the objective is to differentiate between regular and signal functions. (line 2). Discrete events can be modeled as a sum type that indicates whether an event is present or not (line 3). The variant of this type representing a present event can carry additional data (e.g., event timestamp). An event stream denotes a signal of events (line 4). Lastly, the lift functions allow for lifting scalar values to signals (line 5) and regular functions to signal functions by point-wise application (line 6). Equation 1 demonstrates how a lifted operator can be defined using the regular operator.
@@
| 6 | Lifts a function to a signal function | Equation 1 |
@@ 5 Modeling Time Series Data as Signals
The first problem arises when combining asynchronous time series, i.e., the observations have different timestamps. This problem is illustrated in Figure 3 which shows three possibilities to model values between observations: linear interpolation (a), last observation (b), and no value (c). All three graphs contain the same observations, summing up their power consumption based on different models for missing data. The resulting power consumption can vary greatly. For example, the addition yields a piecewise linear function in graph (a), while the addition yields no results in graph (c).
@@
From an FRP position, a time series can be modeled as an event stream from to . This event stream does not assume any value if no observation is present (c). However, based on this event stream, another signal can be created by applying a signal function that can model the values between observations in the event stream. For example, modeling missing values using the last known observation by “holding” the last observation (b) can be expressed using the hold function from [22]. This ability is a powerful tool that allows us to address both of these problems.
@@ 6 A Language-Agnostic Framework
This section introduces a framework of three operators that can be embedded into a knowledge graph query language to support bridging between knowledge graphs and signals. We first present some requirements and explain their importance. Then we formally define the operators that fulfill these requirements. The requirements presented are based on an analysis of the motivating example. The requirements only address the conceptual core of such a query language, not particular computational requirements such as computing the mean of a signal.
@@
Based on these requirements, we introduce a set of operators that form a language-agnostic framework to integrate signals with knowledge graph query languages. If possible, operators from FRP are applied to satisfy the requirements. We use brackets to denote lists.
Signal. The operator constructs a signal based on a non-empty list . For example, evaluates to the signal of . Language-specific implementations of this operator should restrict the elements of to elements of the graph data model. For example, Section 7 restricts the elements to IRIs. This is a knowledge graph-specific operator and has no directly corresponding function in FRP. This operator addresses R1.
@@ 7.1 Data Model
A Signal-Annotated RDF Dataset (SARD) is a triple where is a regular RDF dataset, is a set of RDF signals and is a partial function , where denotes the set of IRIs.
The signal annotation function has two arguments. This is necessary because a single IRI can be attributed with multiple signals. We refer to the first argument of as signal source and the second IRI as signal property. In our example, <Battery1> is a signal source while ev:ActivePower and ev:SoC are signal properties.
@@ 7.3 Semantics
This section describes the semantics of the proposed extension for static knowledge bases. The central idea of the proposed extension is to allow solutions to bind variables to either an RDF term or an RDF signal. For example, in Listing 2, the variable ?charger binds to the IRIs that identify EV chargers, while the variable ?ap binds to a signal that models their active power consumption over time. The semantics of the query language is based on two pillars: the regular SPARQL algebra and FRP. As the RDF dataset within the SARD is used to evaluate regular SPARQL constructs, SigSPARQL does not alter the semantics of, for example, graph pattern matching and filtering solution sequences.
@@ 7.3.2 SIGNALS and WHEN Clause
This is the only means for users to directly create a signal from the RDF dataset. All signal declarations result in the algebraic expression , where is the signal source, is the signal property, and is a new SPARQL algebra expression for the operator from Section 6. For example, obtains the active power signal of the term bound to ?charger. The evaluation will result in an error if the signal source is unbound, is not in , or is undefined. The expression is embedded in a algebra operator222 evaluates an expression and binds the result to a variable in the solution. It is used to realize the BIND(<expr> AS ?var) construct in regular SPARQL. with the corresponding expression and variable name. These operators are inserted right before or immediately after the grouping of solution sequences, depending on whether the entity of the signal is a group condition.
For example, in the motivating example, we may want to sum up all active power sensors in the garage and compare it to the current operating envelope of the garage. This requires grouping the solution sequence by adding a group condition (e.g., GROUP BY ?garage). In this scenario, the operator that uses the AP sensors (not a grouping condition) is inserted before the grouping, while the operator that uses ?garage is inserted after the grouping. Note that none of these variables are within the scope of the GROUP BY and HAVING clauses, even if they are inserted before the grouping.
In addition to the SIGNALS clause, a WHEN clause can be integrated into the SPARQL algebra by creating an operator that wraps the entire query. This therefore runs at the end of the query evaluation process and applies the WHEN clause’s predicate to every solution. Based on the resulting signal, an event stream is created that contains an event when the predicate signal changes from false to true (positive edge). The payload of the events is a regular SPARQL solution that is obtained by evaluating the TSS at the positive edge. If there is a BECOMES TRUE AT statement in the WHEN clause, an additional binding with time information is added to the solution. We will use the term trigger event to refer to the resulting events. See the discussion on the predicate function in Fran [17] for a detailed discussion of this concept.

### v2

@@ Abstract
Purpose: Cyber-Physical Systems (CPSs) integrate computation and physical processes, producing time series data from thousands of sensors. Knowledge graphs can contextualize these data, yet current approaches that are applicably to monitoring CPS rely on observation-based approaches. This limits the ability to express computations on sensor data, especially when no assumptions can be made about sampling synchronicity or sampling rates.
Methodology: We propose an approach for integrating knowledge graphs with signals that model run-time sensor data as functions from time to data. To demonstrate this approach, we introduce SigSPARQL, a query language that can combine RDF data and signals. We assess its technical feasibility with a prototype and demonstrate its use in a typical CPS monitoring use case.
Findings: Our approach enables queries to combine graph-based knowledge with signals, overcoming some key limits of observation-based methods. The developed prototype successfully demonstrated feasibility and applicability.
Value: This work presents a query-based approach for CPS monitoring that integrates knowledge graphs and signals, alleviating problems of observation-based approaches. By leveraging system knowledge, it enables operators to run a single query across different system instances within the same domain. Future work will extend SigSPARQL with additional signal functions and evaluate it in large-scale CPS deployments.
@@
However, fully utilizing sensor data in CPS is challenging. Firstly, data are frequently siloed across various systems. When this occurs, it is essential to have unified data access to effectively combine information from these sources. In addition, even if sensor data access does not pose a problem, some tasks require contextualizing the sensor data. In other words, tasks require information on the CPS itself, not just the pure sensor data. For both aspects, Semantic Web technologies can provide a solution.
One solution is Ontology-Based Data Access (OBDA) [31], which enables the creation of a virtual knowledge graph on-top of specialized data stores (e.g., time series databases). This virtual graph can be queried alongside the context information stored in a graph database, providing contextualized data access without materializing sensor data in the graph. Such an approach has been applied in CPSs use cases (e.g., [34, 5, 24]).
Although OBDA solves many issues in CPSs, computations on time series data can be cumbersome. After mapping a time series to a graph, the temporal values of a single sensor may be spread across thousands of elements that represent individual observations. This can lead to an increase in the complexity of queries, as the concept of a temporal value must be reconstructed from the individual observations.
While efforts have been made to integrate knowledge graphs and time series data (e.g., [7]), they do lack a concept that abstracts over individual observations to alleviate this complexity. As a result, users must resort to “workarounds” to express computations such as summing up the temporal values from two sensors. Although some requirements can be met with these workarounds, there is a mismatch between the conceptual nature of the computation and the formulation in a query language. A Signal [22] is a concept that can abstract over individual observations.
This work proposes to support signals as “first-class citizens” when querying knowledge graphs to overcome the shortcomings of contemporary approaches when formulating computations over sensor data. The integration of this concept involves rethinking the semantics of queries and their results to accommodate the temporal dimension of signals. When considering this integration, the following research questions emerge.
@@
This paper is structured as follows. Section 2 discusses related work, while Section 3 presents a motivating example. Then, Section 4 provides a background on signals, and Section 5 discusses their relationship to time series data. Section 6 presents a list of requirements to bridge the gap between knowledge graphs and signals and proposes a set of operators that address these requirements. Next, Section 7 applies these concepts to the SPARQL query language. Section 8 assesses its feasibility with a prototype, while Section 9 applies the proposed SPARQL extension to the motivating example. Finally, Section 10 reflects on the proposed approach and Section 11 concludes this work.
@@ 2 Related Work
This section will focus on approaches based on Resource Description Framework (RDF), as this work focuses on Semantic Web technologies. Similar approaches may also be applicable to property graphs. We have identified four key areas related to our research topics: (i) RDF Stream Processing, (ii) Querying temporal data with OBDA, (iii) Combining continuous querying and OBDA, and (iv) Integrating graphs with time series data.
RDF Stream Processing: The main objective of RDF Stream Processing (RSP) is to enable continuous queries on streaming graph data. This technology has been used for monitoring CPSs (e.g., [18]). We can broadly distinguish between two RSP approaches: window-based systems and those inspired by Complex Event Processing (CEP).
On the one hand, window-based approaches (e.g., [8, 6, 25, 15]), focus on querying streaming data by dividing unbounded streams into bounded relations using window operators. TEF-SPARQL [17] extends this paradigm by incorporating facts into the query language. Using facts, one can remember assertions beyond a single window. Similar approaches have also been proposed for property graphs [32].
On the other hand, CEP-inspired RSP systems (e.g., [3, 14]) aim to detect patterns in event streams and aggregate them into high-level events. These efforts integrate CEP operations into continuous query frameworks based on SPARQL. OntoEvent [27] proposes formulating CEP queries within a knowledge base. Furthermore, PipeFlow [33], proposes embedding the queries in a data flow programming language, while Teymourian et al. [35] embed SPARQL queries in a logic programming environment.
RSP shares many ideas and properties with temporal RDF graphs [19]. A temporal RDF graph associates a validity interval with each triple. When querying, this validity interval is considered when evaluating a query at a certain time point. RSP systems also annotate their facts with arrival timestamps but focus on continuously updating their results based on new incoming information which is crucial for monitoring CPSs.
Querying temporal data with OBDA: The emergence of the OBDA paradigm, marked by approaches such as Mastro [12] and Ontop [11], has increased the applicability of RDF-based methods in CPSs. This success stems from the concept of a virtual knowledge graph layer built on top of existing databases. Recent CPS-centered approaches focus on building virtual knowledge graphs for time series data. Steindl et al. [34] demonstrated an approach to integrate time series data from OPC UA, a well-known standard in the manufacturing domain. In addition, Chrontext [5] proposes an approach for rewriting SPARQL queries to queries on time series databases. Furthermore, Ontop-temporal [23] extends the querying capabilities with temporal logic formulae.
@@
Approaches integrating graphs with time series data: Recently, the combination of graph and time series data has received increasing attention. Bollen et al. [7] present an approach that extends matching graphs with measurement and time series patterns. The bindings of these patterns can then be used to access the underlying time series. Furthermore, Ammar et al. [2] outline a research roadmap with the goal of creating a hybrid data model for graphs and time series data.
The approach proposed in this work can also be attributed to the latter category. Our approach differentiates itself from all the mentioned works by relying on signals [22] to model temporal values. Using signals allows us to build on existing work created in the Functional Reactive Programming (FRP) community and tackle problems that arise from using an observation-based approach (see Section 5).
@@ 3 A Motivating Example
This section introduces a motivating example of monitoring a CPS that is used throughout this work. Figure 1 shows a conceptual system model of a garage that contains multiple Electric Vehicle (EV) chargers. In addition, the garage includes a photovoltaic system (PV system) and a battery to optimize peak power and overall energy consumption. However, the garage is a substantial power consumer in the area. Therefore, without any further restriction, supporting the garage’s peak power demands would require additional power grid infrastructure, resulting in additional costs for the involved parties.
An envisioned approach to circumvent additional infrastructure is to impose an operating envelope, limiting the maximum power the garage can drain from the grid. This limitation places the responsibility of respecting this threshold on the operator of the EV chargers, who may be held liable for any violations. Thus, the garage operator must be able to detect envelope violations to present alerts to the responsible employees.
@@
Figure 2 shows an extended RDF graph that models a charging garage. In addition to a regular RDF graph (white ovals and solid lines), the extended data model introduces signals (green circles and dashed lines). For now, it suffices to know that the signals model sensor values and that the dashed connections can associate multiple signals with nodes in the RDF graph. These definitions will be detailed in Section 7.1.
In this example, the <Battery1> has two signals. models the active power, while models the battery’s State of Charge (SoC). In addition to the battery, other devices within the garage have an AP signal (e.g., <Charger1>). Lastly, the garage itself has a signal that models the operating envelope. As there is no AP signal at the garage level, we must compute this signal based on the AP of the contained entities (e.g., chargers). Note that the set of contained devices is not known at design-time and must be extracted by querying the knowledge graph. A query that implements this requirement will be presented in Section 9.
@@ 4 Functional Reactive Programming
Assume that the expression SineWave() in Listing 1 creates a signal representing a sine wave. Note that the value of a implicitly depends on time. Therefore, evaluating a at two different time points may yield different results. As a is a signal, the expression b = a * 2 denotes that b is twice the value of a at any given time point. This is different from denoting that b is twice the current value of a in imperative programming languages. In FRP, if a changes its value, the signal associated with b automatically reflects this change, allowing b and c to become scaled sine waves. The concept of a signal is very flexible and allows us to model sensor values over time.
Table 1 contains descriptions of relevant FRP concepts and relates these concepts to the motivating example. Formally, a signal (row 1) is a partial function of a time domain to a value domain . The value of a signal at a time point is denoted by . We refer to any function that has signals as input or output as signal function111This is a more lenient usage than the original definition [22], which restricts signal functions to achieve favorable computational properties. In this work, the objective is to differentiate between regular and signal functions. (row 2). Discrete events can be modeled as a sum type that indicates whether an event is present or not (row 3). The variant of this type representing a present event can carry additional data (e.g., event timestamp). An event stream denotes a signal of events (row 4). Lastly, the lift functions allow for lifting scalar values to signals (row 5) and regular functions to signal functions by point-wise application (row 6).
@@
| 6 | Lifts a function to a signal function |
@@ 5 Modeling Time Series Data as Signals
The first challenge arises when combining asynchronous time series, where observations have different timestamps. Figure 3 shows the addition of two power sensors and illustrates three ways to model values between observations: linear interpolation (a), last observation (b), and no value (c). Although all three graphs use the same observations, the observed results differ significantly.
@@
From an FRP position, a time series can be modeled as an event stream from to that is analogous to (c). However, based on this event stream, another signal can be created by applying a signal function that models the values between observations. For example, modeling missing values using the last known observation by “holding” the last observation (b) can be expressed using the hold function from [22]. This ability is a powerful tool that allows us to address both of these problems.
@@ 6 A Language-Agnostic Framework
This section introduces a framework of three operators that can be embedded in a knowledge graph query language to support the integration of knowledge graphs and signals. We first present some requirements and explain their importance. Then we define operators that fulfill these requirements. The requirements presented are based on an analysis of the motivating example. They address the conceptual core of such a query language, not particular computational requirements such as computing the mean of a signal.
@@
Based on these requirements, we introduce a set of operators that form a language-agnostic framework to integrate signals with knowledge graph query languages. If possible, operators from FRP are applied to satisfy the requirements.
Signal. The operator constructs a signal based on a signal source and a signal property . For example, evaluates to the signal of . Language-specific implementations of this operator can restrict the elements of and to elements of the used data model. For example, Section 7 restricts the elements to IRIs. This operator addresses R1.
@@ 7.1 Data Model
A signal-annotated RDF dataset is a triple where is a regular RDF dataset, is a set of RDF signals, and is a partial function , where denotes the set of IRIs.
The signal annotation function has two arguments. This is necessary because a single IRI can be attributed with multiple signals (e.g., <Battery1> from Figure 2). We refer to the first IRI of as signal source and the second IRI as signal property. In our example, <Battery1> is a signal source while ev:ActivePower and ev:SoC are signal properties.
@@ 7.3 Semantics
This section describes the semantics of the proposed extension for static knowledge bases. The central idea of the proposed extension is to allow solutions to bind variables to either an RDF term or an RDF signal. For example, in Listing 2, the variable ?charger binds to the IRIs that identify EV chargers, while the variable ?ap binds to a signal that models their active power consumption over time. The semantics of the query language is based on two pillars: the regular SPARQL algebra and FRP. The regular RDF dataset, which is part of the signal-annotated RDF dataset, is utilized for the evaluation of regular SPARQL constructs. Consequently, SigSPARQL does not modify the semantics of graph pattern matching and filtering solution sequences.
@@ 7.3.2 SIGNALS and WHEN Clause
The SIGNALS clause is the only means for users to directly create a signal. All signal declarations result in the algebraic expression , where is the signal source, is the signal property, and is a new expression that implements the operator from Section 6. For example, obtains the AP signal of the term bound to ?garage. Evaluating for a solution will result in an error if is unbound, is not an IRI, or if is undefined.
The expression is embedded in an graph pattern from the SPARQL algebra [21]. can be used to evaluate an expression and bind the result to a new variable for every solution. For example, the SPARQL construct BIND(?a * ?b AS ?c) computes the value of ?a * ?b and binds it to the variable ?c. The resulting graph patterns are then inserted right before or immediately after the grouping of solution sequences, depending on whether the signal source is a group condition.
For example, in the motivating example, one must sum up all active power sensors in the garage and compare it to the current operating envelope of the garage. This requires grouping the solution sequence by adding a group condition (e.g., GROUP BY ?garage). In this scenario, the graph pattern that uses the AP sensors (not a grouping condition) is inserted before the grouping, while the graph pattern that uses ?garage is inserted after the grouping. Note that none of these variables are within the scope of the GROUP BY and HAVING clauses, even if they are inserted before the grouping.
In addition to the SIGNALS clause, a WHEN clause can be integrated into the SPARQL algebra by creating an graph pattern that wraps the entire query. This therefore runs at the end of the query evaluation process and applies the WHEN clause’s predicate to every solution. Based on the resulting signal, an event stream is created that contains an event when the predicate signal changes from false to true (positive edge). In the motivating example, the positive edge denotes every time point where the envelope becomes violated. The payload of the events is a regular SPARQL solution that is obtained by evaluating the TSS at the positive edge. If there is a BECOMES TRUE AT statement in the WHEN clause, an additional binding with time information is added to the solution. We will use the term trigger event to refer to the resulting events. See the discussion on the predicate function in Fran [16] for a detailed discussion of this concept.